# Envelopes in the class of Banach algebras of polynomial growth and $C^{\infty}$-functions of a finite number of free variables

### 11 Jan 2024 v1

@@ Abstract.
The Arens–Michael envelope can be treated a functor from the algebraic geometry to the analytic geometry (at the level of function algebras). In particular, it corresponds to the analytization of an affine scheme of finite type; see [35]. An advantage of this view on the analytization is that the functor is well defined for non-commutative algebras as well. Although this approach also has its drawbacks (for example, algebras with rich representation theory such as the Weyl algebra can vanish), it works well in many cases; see, e.g. [4]. We are interested in finding a similar functor from the algebraic geometry to the differential geometry, that is, from the category of associative real algebras to some category of real topological associative algebras that exhibits the same behaviour as the Arens–Michael envelope in the analytic case. This means that it must take the algebra of real polynomials $\mathbb{R}[x_{1},\ldots,x_{k}]$, $k\in\mathbb{N}$, into the algebra $C^{\infty}(\mathbb{R}^{k})$ of real-valued infinitely differentiable functions. Moreover, it is natural to expect that the functor should act non-trivially on a number of classical examples of non-commutative algebras.
@@
The axiomatic approach based on the class of Banach algebras of polynomial growth is indeed adequate to the problem since the natural embedding of $\mathbb{R}[x_{1},\ldots,x_{m}]\to C^{\infty}(\mathbb{R}^{m})$ turns out to be an envelope with respect to this class; see Proposition 3.3. By definition, the envelope is an algebra that is locally of polynomial growth (contained in the class $\mathsf{PGL}$ of projective limits algebras of polynomial growth) and is universal in the sense that any homomorphism into a Banach algebra of polynomial growth factors through it. The envelope exists for each real associative algebra $A$ and, furthermore, for each real topological algebra. In fact, first examples were considered by the author in [5], although the corresponding terminology is not introduced there.
@@
Note that the algebras of ‘smooth functions’ traditionally considered in the non-commutative differential geometry in the spirit of Connes are usually self-adjoint subalgebras of $C^{*}$-algebras (in particular, they have trivial radical). But, in our reasoning, algebras of triangular matrices is fundamental and, as a consequence, the examples discussed in this paper admit a (trivial) involution only in the commutative case. This restriction appears to be an essential feature of our approach. Furthermore, the presented theory has the shortcoming that it does not include some classical examples, such as the non-commutative tori (since complex numbers are used in their definition) and the quantum disc (see Example 4.1).
@@
Although theoretical constructions occupy a significant part of this text, nevertheless, the main focus is on examples. In addition to free algebras, we consider the universal enveloping algebras of finite-dimensional Lie algebras (refining results in [5]), the coordinate algebras of the quantum plane and quantum group $SL_{q}(2,\mathbb{R})$; see § 4. The commutative case is also of interest and discussed in § 3. We discuss the general properties of the class $\mathsf{PGL}$ in § 2 and also give a universal construction of $C^{\infty}$-tensor algebra. Note that the $C^{\infty}$-tensor algebra of a $k$-dimensional space is exactly the algebra of ‘free $C^{\infty}$-functions’ of rank $k$.
@@
Note that (in the case of real algebras) the second and third classes both contain $\mathsf{PG}$, which plays a main role in this article. The envelopes with respect to the third class is introduced in [7] and also some examples are given there. The stability properties for strictly real Banach algebras are proved in [31, Theorem 4.4] but the envelopes have not been studied yet. Note also that for algebras with involution, envelopes with respect to the class of $C^{*}$-algebras were studied earlier; see, for example, [19, 27].
The author wishes to thank the Institute of Mathematics and Mechanics at the Kazan Federal University and especially to Renat Gumerov for the hospitality during the visit to Kazan in 2023.
@@ 1. Local containment and envelopes
We denote by $\mathsf{TA}$ the category of topological algebras $\mathbb{C}$ or $\mathbb{R}$, that is, topological vector spaces endowed with separately continuous multiplication; see [30, Chapter 1, § 1, p. 4, Definition 1.1] or [20, § 1, p. 6, Definition 1.6]. In general, the underlying vector space is not assumed locally convex or complete and the existence of identity is also not assumed. Here the morphisms are continuous homomorphisms. Note that a subclass of objects in $\mathsf{TA}$ can be treated as a full subcategory. Following the tradition of functional analysis, we use the slightly old-fashioned term projective limit (aka ‘inverse limit’) for what is called directed limit in category theory.
@@
Recall that a topological algebra whose topology can be determined by a family of submultiplicative seminorms is called an Arens–Michael algebra. If $\mathsf{C}$ is contained in the class of all Banach algebras, then each algebra locally contained in $\mathsf{C}$ is obviously an Arens–Michael algebra. Furthermore, an Arens–Michael algebra is exactly an algebra that is locally in the class of all Banach algebras. Note also that the limit of a projective system of Banach algebras in $\mathsf{TA}$ coincides with the limit in the category of Arens–Michael algebras; see, e.g., [30, Chapter 3, § 2, p. 84, (2.8)].
@@
Of course, the definition makes sense not only for $\mathsf{CL}$ but also for an arbitrary full subcategory in $\mathsf{TA}$, but we do not need it here. Moreover, in the general situation it natural to modify the definition; see [8].
The choice of the term ‘envelope’ is more likely due to historical considerations; cf. [21, Chapter 5]. It might as well be called ’free functor’. However, we reserve the last term for those cases where the forgetful functor changes the algebraic type of the category (such as ‘associative algebras $\rightarrow$ vector spaces’ or ‘associative algebras $\rightarrow$ sets’). Akbarov in his papers [2, 3] also uses the term ‘envelope’, which has a similar but different meaning.
@@
The question of sufficient conditions for the existence of an envelope (and a left adjoint to the forgetful functor) can be solved quite simply. This can be seen from the following proposition.
@@
Given a topological algebra $A$, consider the set of all continuous submultiplicative seminorms whose completions belong to $\mathsf{C}$ and the corresponding family of continuous homomorphisms from $A$ to the completions. Taking the closures of the images of $A$ in all possible finite products, it is easy to obtain linking homomorphisms that form a directed system in $\mathsf{C}$. Its limit in $\mathsf{TA}$ obviously satisfies the conditions in the definition of the envelope. ∎
@@
Of course, to verify that the tensor $\mathsf{CL}$ algebra exists for each $E$ we need to prove the existence of a left adjoint functor. To do this we use analytic tensor algebras.
If $\mathsf{B}$ is the class of all Banach algebras, then the tensor $\mathsf{BL}$ algebra is called the analytic tensor algebra (or Arens-Michael tensor algebra) and denoted by $\widehat{T}(E)$. In the case of the ground field $\mathbb{C}$, see a proof of the existence and an explicit construction in [34, § 4.2]. For $\mathbb{R}$ the proof is identical.
@@ 2. Banach algebras of polynomial growth and $\mathsf{PGL}$ algebras
We consider $\mathsf{PGL}$ algebras, that is, $\mathsf{CL}$ algebras in the case when $\mathsf{C}$ is the class of Banach algebras of polynomial growth.
@@ 3. Envelopes of certain commutative algebras
Envelopes with respect to the class $\mathsf{PG}$
@@
The natural embedding $\mathbb{R}[x_{1},\ldots,x_{k}]\to C^{\infty}(\mathbb{R}^{k})$, $k\in\mathbb{N}$, is an envelope with respect to $\mathsf{PG}$.
@@
A proof of a similar result for complex-valued functions of class $C^{\infty}$ and a complex Banach algebra can be found, for example, in [10, Theorem 1]. In the case of a real Banach algebra, it suffices to apply the result to the complexification of the algebra $B$ and note that $\Phi_{V}$ sends all the real-valued functions to $B$. The latter follows from the fact that $\mathbb{R}[x]$ is dense in $C^{\infty}(V)$. ∎
@@
(B) Every $C^{\infty}(\mathbb{R}^{m})$-submodule of the module $C^{\infty}(\mathbb{R}^{k})^{p}$, $p\in\mathbb{N}$, generated by a finite set of $p$-tuples of polynomials (or even of real analytic functions) is closed.
It is obvious that Part (A) follows from Part (B). For a proof of Part (B) see [39, Chapitre VI, p. 119, Corollaire 1.5]. See also a proof of Part (A) in [29, Théorème 4] and partial cases in [23, 28].
@@
determines a continuous linear map $\widehat{\varphi}$ that is a homomorphism since $XY=0$. Thus $\widehat{\varphi}$ is an desired extension of $\varphi$. Finally, it is easy to see that $\mathcal{R}(\Omega)$ is dense in $C^{\infty}(\Omega)$. Therefore $\widehat{\varphi}$ is unique.
Example 3.8.
Consider the algebra $A=\mathbb{R}[x,y]/(x^{2}-y^{3})$. Using the parametrization $x=t^{3}$, $y=t^{2}$, we can identify $A$ with the subalgebra $P_{1}$ in $\mathbb{R}[t]$ consisting of polynomials with zero coefficient at the term of first degree, i.e., with trivial first derivative at 00. We now show that the same holds for the envelope, i.e., $\widehat{A}^{\,\mathsf{PG}}\cong C^{\infty}_{1}$, where
$C^{\infty}_{1}\!:=\{f\in C^{\infty}(\mathbb{R})\!:\,f^{\prime}(0)=0\}.$
We claim first that if $f\in C^{\infty}(\mathbb{R})$ and even, then the function $g(y)\!:=f(y^{1/2})$ is in $C^{\infty}[0,+\infty)$. Indeed, it is obvious that $g$ is of class $C^{\infty}$ for $y>0$. In the case when $y=0$, take $n\in\mathbb{N}$ and write $f(t)=P_{2n}(t)+o(t^{2n})$ as $t\to 0$, where $P_{k}$ denotes the Taylor polynomial of degree $k$ at $t=0$. Since $P_{2n}(y^{1/2})$ is a polynomial in $y$ and $g(y)-P_{2n}(y^{1/2})=o(y^{n})$ as $y\to 0+$, we have that $g$ is $n$ times differentiable at 00. Since $n$ is arbitrary, $g\in C^{\infty}[0,+\infty)$.
We claim second that if $f$ is an odd function in $C^{\infty}(\mathbb{R})$ and $f^{\prime}(0)=0$, then the function $h(y)\!:=f(y^{1/2})/y^{3/2}$ extends to a function in $C^{\infty}[0,+\infty)$. Indeed, it is easy to see from Taylor’s formula, that the function $t\mapsto f(t)/t^{3}$ extends to a function in $C^{\infty}(\mathbb{R})$ and so we can apply the first claim.
Let $B$ be a Banach algebra of polynomial growth and $\varphi\!:P_{1}\to B$ a homomorphism. Denote by $\Psi$ the $C^{\infty}$-functional calculus $C^{\infty}(\mathbb{R})\to B$ associated with the element $\varphi(t^{2})$. Write $f=f_{+}+f_{-}$, where $f_{+}$ is even, $f_{-}$ is odd and both functions are in $C^{\infty}(\mathbb{R})$. Put $g(y)\!:=f_{+}(y^{1/2})$, $h(y)\!:=f_{-}(y^{1/2})/y^{3/2}$ and
$\widehat{\varphi}\!:C^{\infty}_{1}\to B\!:f\mapsto\Psi(g)+\Psi(h)\varphi(t^{3}).$
The maps $f\mapsto f_{+}$ and $f\mapsto f_{-}$ are obviously continuous and so is $f\mapsto\tilde{f}$ since $f\mapsto f\,t^{3}$ has closed range by Proposition 3.6. Therefore $\widehat{\varphi}$ is also continuous.
If $p\in P_{1}$, then evidently $p=p_{+}+\tilde{p}\,t^{3}$ and hence $\widehat{\varphi}(p)=\varphi(p)$. Therefore $\widehat{\varphi}$ extends $\varphi$. It is easy to see that $P_{1}$ is dense in $C^{\infty}_{1}$. Hence $\widehat{\varphi}$ is a homomorphism and, moreover, a unique homomorphism with the desired properties. Thus $\widehat{C}^{\,\mathsf{PG}}\cong C^{\infty}_{1}$.
@@ 4. Envelopes of certain non-commutative algebras
(A) The multiplication determined by the relations $xy=qyx$ and $xy=u$ extends to a continuous operation in $C^{\infty}(\Omega)\mathbin{\widehat{\otimes}}\mathbb{R}[[u]]$ that turns it into a $\mathsf{PGL}$ algebra.
@@
It follows that the upper right entries of $\widetilde{\pi}_{p,\lambda}(u)^{r}$ and $\widetilde{\pi}^{\prime}_{p,\mu}(u)^{r}$ vanish when $r\neq p$. Therefore $\rho^{\prime}$ is the product of maps $\rho^{\prime}_{r}\!:C^{\infty}(\Omega)\to C^{\infty}(\mathbb{R})\times C^{\infty}(\mathbb{R})$ (where $r=p-1$). Write elements as in (4.1) and note that $\widetilde{\pi}_{r+1,\lambda}(f_{r})$ is the diagonal matrix with the entries $f_{r}(\lambda q^{r}),\ldots,f_{r}(\lambda)$ and $\widetilde{\pi}_{r+1,\lambda}(g_{r})$ has $g_{r}(0)$ at each diagonal entry. Applying the equality $f_{r}(0)=g_{r}(0)$, we conclude that for every $r$,
@@
So $\rho^{\prime}_{r}$ is the restriction to $C^{\infty}(\Omega)$ of the product of two continuous linear endomorphism of $C^{\infty}(\mathbb{R})$ each of which is a composition of three linear maps, the first is induced by a diffeomorphism, the second is the multiplication operator on a polynomial and the third is a shift by a one-dimension operator. Every of these maps has closed range, the second map by Proposition 3.6 and the first and third because they are topological isomorphisms. Also, it is easy to see that all the maps are injective. Thus, by the inverse mapping theorem for Fréchet spaces, they are topologically injective and so is $\rho^{\prime}$ by Part (A) of Lemma 4.4. This completes the proof of Part (A).
(B) Let $B$ be a Banach algebra of polynomial growth and $\varphi\!:\mathcal{R}(\mathbb{R}^{2}_{q})\to B$ a homomorphism. Denote by $\theta_{x}$ and $\theta_{y}$ the $C^{\infty}$-functional calculi (i.e., continuous homomorphisms from $C^{\infty}(\mathbb{R})$ to $B$) corresponding to $\varphi(x)$ and $\varphi(y)$. Recall that $C^{\infty}(\Omega)=\{(f,g)\in C^{\infty}(\mathbb{R})\times C^{\infty}(\mathbb{R})\!:\,f(0)=g(0)\}$ by definition. Consider the continuous linear map
@@
Note that $\varphi(xy)$ is nilpotent by Lemma 4.2. Then $\mathbb{R}[u]\to B\!:u\mapsto\varphi(xy)$ can be extended to continuous linear map $\mathbb{R}[[u]]\to B$ (this a partial case of Theorem 4.5). Take the composition of the tensor product of these two maps and the linearization of the multiplication in $B$,
@@
(A) By Theorem 2.4, every commutator belongs to the radical. In particular, $bc=(q-q^{-1})^{-1}[a,d]$ and so belongs to the radical; cf. Lemma 4.2. Then $bc$ is nilpotent again by Theorem 2.4. Hence $1+q^{-1}bc$ and $1+qbc$ are invertible. It follows from (4.7) that $a$ and $d$ are also invertible.
@@
is a PBW-basis in $A_{q}$. Thus we immediately obtain the following result.
@@
(A) The multiplication in $A_{q}$ determined by the relations in (4.5) extends to a continuous operation in $C^{\infty}(SL_{q}(2,\mathbb{R}))$ that turns it into a $\mathsf{PGL}$ algebra.
@@
(A) The multiplication in $U(\mathfrak{g})$ extends to a continuous operation in $C^{\infty}_{\mathfrak{g}}$, which makes it a $\mathsf{PGL}$ algebra,
@@
Recall that $\mathcal{R}(SL_{q}(2,\mathbb{R}))$ and $U(\mathfrak{g})$ have Hopf algebra structures. Similarly, $\widehat{U}(\mathfrak{g})^{\,\mathsf{PG}}$ and ${\mathcal{R}(SL_{q}(2,\mathbb{R}))^{^}}{}^{\,\mathsf{PG}}$ admit induced topological Hopf algebra structures. This topic is discussed in [8].

### 11 Mar 2024 v2

@@ Abstract.
The Arens–Michael envelope can be treated a functor from the algebraic geometry to the analytic geometry (at the level of function algebras). In particular, it corresponds to the analytization of an affine scheme of finite type; see [34]. An advantage of this view on the analytization is that the functor is well defined for non-commutative algebras as well. Although this approach also has its drawbacks (for example, algebras with rich representation theory such as the Weyl algebra can vanish), it works well in many cases; see, e.g., [4]. We are interested in finding a similar functor from the algebraic geometry to the differential geometry, that is, from the category of associative real algebras to some category of real topological associative algebras that exhibits the same behaviour as the Arens–Michael envelope in the analytic case. This means that it must take the algebra of real polynomials $\mathbb{R}[x_{1},\ldots,x_{k}]$, $k\in\mathbb{N}$, into the algebra $C^{\infty}(\mathbb{R}^{k})$ of real-valued infinitely differentiable functions. Moreover, it is natural to expect that the functor should act non-trivially on a number of classical examples of non-commutative algebras.
@@
The axiomatic approach based on the class of Banach algebras of polynomial growth is indeed adequate to the problem since the natural embedding of $\mathbb{R}[x_{1},\ldots,x_{m}]\to C^{\infty}(\mathbb{R}^{m})$ turns out to be an envelope with respect to this class; see Proposition 3.3. By definition, the envelope is an algebra that is locally of polynomial growth (contained in the class $\mathsf{PGL}$ of projective limits algebras of polynomial growth) and is universal in the sense that any homomorphism into a Banach algebra of polynomial growth factors through it. The envelope exists for each real associative algebra $A$ and, furthermore, for each real topological algebra. In fact, first examples were considered by the author in [5], although the corresponding terminology was not introduced there.
@@
Note that the algebras of ‘smooth functions’ traditionally considered in the non-commutative differential geometry in the spirit of Connes are usually self-adjoint subalgebras of $C^{*}$-algebras (in particular, they have trivial radical). But, in our reasoning, algebras of triangular matrices are fundamental and, as a consequence, the examples discussed in this paper admit a (trivial) involution only in the commutative case. This restriction appears to be an essential feature of our approach. Furthermore, the presented theory has the shortcoming that it does not include some classical examples, such as the non-commutative tori (since complex numbers are used in their definition) and the quantum disc (see Example 4.1).
@@
Although theoretical constructions occupy a significant part of this text, nevertheless, the main focus is on examples. In addition to free algebras, we consider the universal enveloping algebras of finite-dimensional Lie algebras (refining results in [5]), the coordinate algebras of the quantum plane and quantum group $SL_{q}(2,\mathbb{R})$; see § 4. The commutative case is also of interest and discussed in § 3. We discuss the general properties of the class $\mathsf{PGL}$ in § 2 and also give a universal construction of $C^{\infty}$-tensor algebra in § 1. Note that the $C^{\infty}$-tensor algebra of a $k$-dimensional space is exactly the algebra of ‘free $C^{\infty}$-functions’ of rank $k$.
@@
Note that (in the case of real algebras) the second and third classes both contain $\mathsf{PG}$, which plays a main role in this article. The envelopes with respect to the third class is introduced in [7] and also some examples are given there. The stability properties for strictly real Banach algebras are proved in [30, Theorem 4.4] but the envelopes have not been studied yet. Note also that for algebras with involution, envelopes with respect to the class of $C^{*}$-algebras were studied earlier; see, for example, [18, 26].
The author wishes to thank the Institute of Mathematics and Mechanics at the Kazan Federal University and especially Renat Gumerov for the hospitality during the visit to Kazan in 2023.
@@ 1. Local containment and envelopes
We denote by $\mathsf{TA}$ the category of topological algebras over $\mathbb{C}$ or $\mathbb{R}$, that is, topological vector spaces endowed with separately continuous multiplication; see [29, Chapter 1, § 1, p. 4, Definition 1.1] or [19, § 1, p. 6, Definition 1.6]. In general, the underlying vector space is not assumed locally convex or complete and the existence of identity is also not assumed. Here the morphisms are continuous homomorphisms. Note that a subclass of objects in $\mathsf{TA}$ can be treated as a full subcategory. Following the tradition of functional analysis, we use the slightly old-fashioned term projective limit (aka ‘inverse limit’) for what is called directed limit in category theory.
@@
Recall that a topological algebra whose topology can be determined by a family of submultiplicative seminorms is called an Arens–Michael algebra. If $\mathsf{C}$ is contained in the class of all Banach algebras, then each algebra locally contained in $\mathsf{C}$ is obviously an Arens–Michael algebra. Furthermore, an Arens–Michael algebra is exactly an algebra that is locally in the class of all Banach algebras. Note also that the limit of a projective system of Banach algebras in $\mathsf{TA}$ coincides with the limit in the category of Arens–Michael algebras; see, e.g., [29, Chapter 3, § 2, p. 84, (2.8)].
@@
Of course, the definition makes sense not only for $\mathsf{CL}$ but also for an arbitrary full subcategory in $\mathsf{TA}$, but we do not need it here.
The choice of the term ‘envelope’ is more likely due to historical considerations; cf. [20, Chapter 5]. It might as well be called ’free functor’. However, we reserve the last term for those cases where the forgetful functor changes the algebraic type of the category (such as ‘associative algebras $\rightarrow$ vector spaces’ or ‘associative algebras $\rightarrow$ sets’). Akbarov in his papers [2, 3] also uses the term ‘envelope’, which has a similar but different meaning.
@@
The following proposition gives sufficient conditions for the existence of an envelope.
@@
Given a topological algebra $A$, consider the set of all continuous submultiplicative seminorms on $A$ whose completions belong to $\mathsf{C}$ and the corresponding family of continuous homomorphisms from $A$ to the completions. Taking the closures of the images of $A$ in all possible finite products, it is easy to obtain linking homomorphisms that form a directed system in $\mathsf{C}$. Its limit in $\mathsf{TA}$ obviously satisfies the conditions in the definition of the envelope. ∎
@@
Of course, to verify that a tensor $\mathsf{CL}$ algebra exists for each $E$ we need to prove the existence of a left adjoint functor. To do this we use analytic tensor algebras.
If $\mathsf{B}$ is the class of all Banach algebras, then the tensor $\mathsf{BL}$ algebra is called the analytic tensor algebra (or Arens-Michael tensor algebra) and denoted by $\widehat{T}(E)$. In the case of the ground field $\mathbb{C}$, see a proof of the existence and an explicit construction in [33, § 4.2]. For $\mathbb{R}$ the proof is identical.
@@ 2. Banach algebras of polynomial growth and $\mathsf{PGL}$ algebras
In this section, we consider $\mathsf{PGL}$ algebras, that is, $\mathsf{CL}$ algebras in the case when $\mathsf{C}$ is the class of Banach algebras of polynomial growth.
@@ 3. Envelopes of certain commutative algebras
Envelopes with respect to $\mathsf{PG}$
@@
The natural embedding $\mathbb{R}[x_{1},\ldots,x_{n}]\to C^{\infty}(\mathbb{R}^{n})$, $n\in\mathbb{N}$, is an envelope with respect to $\mathsf{PG}$.
@@
A proof of a similar result for complex-valued functions of class $C^{\infty}$ and a complex Banach algebra can be found, for example, in [9, Theorem 1]. In the case of a real Banach algebra, it suffices to apply the result to the complexification of the algebra $B$ and note that $\Phi_{V}$ sends all the real-valued functions to $B$. The latter follows from the fact that $\mathbb{R}[x]$ is dense in $C^{\infty}(V)$. ∎
@@
(B) Every $C^{\infty}(\mathbb{R}^{k})$-submodule of the module $C^{\infty}(\mathbb{R}^{k})^{p}$, $p\in\mathbb{N}$, generated by a finite set of $p$-tuples of polynomials (or even of real analytic functions) is closed.
It is obvious that Part (A) follows from Part (B). For a proof of Part (B) see [38, Chapitre VI, p. 119, Corollaire 1.5]. See also a proof of Part (A) in [28, Théorème 4] and partial cases in [22, 27].
@@
determines a continuous linear map $\widehat{\varphi}$ that is a homomorphism since $XY=0$. Thus $\widehat{\varphi}$ is a desired extension of $\varphi$. Finally, it is easy to see that $\mathcal{R}(\Omega)$ is dense in $C^{\infty}(\Omega)$. Therefore $\widehat{\varphi}$ is unique.
@@ 4. Envelopes of certain non-commutative algebras
(A) The multiplication determined by the relations $xy=qyx$ and $xy=u$ extends to a continuous operation on $C^{\infty}(\Omega)\mathbin{\widehat{\otimes}}\mathbb{R}[[u]]$ that turns it into a $\mathsf{PGL}$ algebra.
@@
It follows that the upper right entries of $\widetilde{\pi}_{p,\lambda}(u)^{r}$ and $\widetilde{\pi}^{\prime}_{p,\mu}(u)^{r}$ vanish when $r\neq p$. Putting $r=p-1$, we can write $\rho^{\prime}$ as the product of maps $\rho^{\prime}_{r}\!:C^{\infty}(\Omega)\to C^{\infty}(\mathbb{R})\times C^{\infty}(\mathbb{R})$. The explicit form of $\rho^{\prime}_{r}$ is as follows. Let elements be written as in (4.1) and note that $\widetilde{\pi}_{r+1,\lambda}(f_{r})$ is the diagonal matrix with the entries $f_{r}(\lambda q^{r}),\ldots,f_{r}(\lambda)$ and $\widetilde{\pi}_{r+1,\lambda}(g_{r})$ has $g_{r}(0)$ at each diagonal entry. Applying the equality $f_{r}(0)=g_{r}(0)$, we conclude that for every $r$,
@@
So $\rho^{\prime}_{r}$ is the restriction to $C^{\infty}(\Omega)$ of the product of two continuous linear endomorphism of $C^{\infty}(\mathbb{R})$ each of which is a composition of three linear maps, the first is induced by a diffeomorphism, the second is the multiplication operator on a polynomial and the third is a shift by a one-dimension operator. Each of these maps has closed range, the second map by Proposition 3.6 and the first and third because they are topological isomorphisms. Also, it is easy to see that all three maps are injective. Thus, by the inverse mapping theorem for Fréchet spaces, they are topologically injective and so is $\rho^{\prime}$ by Part (A) of Lemma 4.4. This completes the proof of Part (A).
(B) Suppose now that $q\neq 1$. Let $B$ be a Banach algebra of polynomial growth and $\varphi\!:\mathcal{R}(\mathbb{R}^{2}_{q})\to B$ a homomorphism. Denote by $\theta_{x}$ and $\theta_{y}$ the $C^{\infty}$-functional calculi (i.e., continuous homomorphisms from $C^{\infty}(\mathbb{R})$ to $B$) corresponding to $\varphi(x)$ and $\varphi(y)$. Recall that $C^{\infty}(\Omega)=\{(f,g)\in C^{\infty}(\mathbb{R})\times C^{\infty}(\mathbb{R})\!:\,f(0)=g(0)\}$ by definition. Consider the continuous linear map
@@
Note that $\varphi(u)$ is nilpotent by Lemma 4.2. Then $\mathbb{R}[u]\to B\!:u\mapsto\varphi(u)$ can be extended to continuous linear map $\mathbb{R}[[u]]\to B$ (this a partial case of Theorem 4.5). Take the composition of the tensor product of these two maps and the linearization of the multiplication in $B$,
@@
(A) By Theorem 2.4, every commutator belongs to the radical. In particular, $bc=(q-q^{-1})^{-1}[a,d]$ and so belongs to the radical; cf. Lemma 4.2. Then $bc$ is nilpotent by Theorem 2.4. Hence $1+q^{-1}bc$ and $1+qbc$ are invertible. It follows from (4.7) that $a$ and $d$ are also invertible.
@@
is a PBW basis in $A_{q}$. Thus we immediately obtain the following result.
@@
(A) The multiplication in $A_{q}$ determined by the relations in (4.5) extends to a continuous operation on $C^{\infty}(SL_{q}(2,\mathbb{R}))$ that turns it into a $\mathsf{PGL}$ algebra.
@@
(A) The multiplication in $U(\mathfrak{g})$ extends to a continuous operation on $C^{\infty}_{\mathfrak{g}}$, which makes it a $\mathsf{PGL}$ algebra,
@@
Recall that $\mathcal{R}(SL_{q}(2,\mathbb{R}))$ and $U(\mathfrak{g})$ have Hopf algebra structures. Similarly, $\widehat{U}(\mathfrak{g})^{\,\mathsf{PG}}$ and ${\mathcal{R}(SL_{q}(2,\mathbb{R}))^{^}}{}^{\,\mathsf{PG}}$ admit induced topological Hopf algebra structures. This topic will be discuss in a separate paper.